A 45.0 mL sample of 0.0100 M Cu2+ buffered at pH 8.00 is titrated with 0.0250 M EDTA. Calculate pCu2+ at each of the added volumes of EDTA. For the Cu2+−EDTA complex, log⁡Kf is 18.78. The fraction of free EDTA in the Y4− form, αY4−, can be found in the table below.

0 mLpCu2+=

6.00 mLpCu2+=

18.0 mLpCu2+=

23.0 mLpCu2+=

pH αY4-
0 1.3 x 10-23
1 1.4 x 10-18
2 2.6 x 10-14
3 2.1 x 10-11
4 3.0 x 10-9
5 2.9 x 10-7
6 1.8 x 10-5
7 3.8 x 10-4
8 4.2 x 10-3
9 0.041
10 0.30
11 0.81
12 0.98
13 1.00
14 1.00
Answer #1

Volume of Cut2 solution = 45.0 mL Conc. of Cu+2 solution = 0.0100 M Conc.of EDTA = 0.0250 M logKf = 18.78 Formation constant

c) Addition of 18.0 mL of EDTA: mmols of Cu+2 = M*V = 0.01 M* 15.0 mL = 0.45 mmol mmols of EDTA = M*V = 0.0250 * 18.0 mL = 0.

0.007143 . (: x is small) CuY-21 [Cu ][Y-+] 2 0.007143 0.007143 - X X.X 0.007143 3x = -K 2. 82x10-19 2.53 x 1016 x=5.31x10-10

K = [Cuy?) -=1.42 x 10-16 ^4 *[Cu*?][Y+] 10+27 [CuY-?] =[Cu*] = 0.006618 Kyx[Y+] 2.53x106 x 0.001838 [Cu*?]=1.42x10-16 = pCu